# Efficiently Vectorized MCMC on Modern Accelerators

### v1

@@ 1   Introduction
Automatic vectorization is the act of transforming one function into another that can handle batches of inputs without user intervention. Implementations of automatic vectorization algorithms—such as JAX’s vmap—are now available in many mainstream scientific computing libraries, and have dramatically simplified the task of running multiple instances of a single algorithm concurrently. They are routinely used to train neural networks (Flax,, 2023) and in other scientific applications, e.g., Schoenholz and Cubuk, (2021); Oktay et al., (2023); Pfau et al., (2020).
This paper focuses on the use of automatic vectorization for Markov chain Monte Carlo (MCMC) algorithms. Tools like vmap provide a convenient way to run multiple MCMC chains in parallel: one can simply write single-chain MCMC code, and call vmap to turn it into vectorized, multi-chain code that can run in parallel on the same processor. Many state-of-the-art MCMC libraries have consequently adopted machine learning frameworks with automatic vectorizaton tools—such as JAX or TensorFlow—as their backend.
One limitation with automatic vectorization tools in modern frameworks is how they handle control flow. Since all instructions must be executed in lock-step, if the algorithm has a while loop all chains must wait until the last chain has finished its iterations. This can lead to serious inefficiencies for MCMC algorithms that generate each sample using variable-length while loops. Roughly speaking, if vectorization executes, say, 100 chains in parallel, all but one finish after at most 10 steps, and the remaining chain runs for 1000 steps, then about 99% of the GPU capacity assigned to vmap is wasted (and our simulations show that the effect can indeed be this drastic). For the No-U-Turn Sampler (HMC-NUTS) (Hoffman et al.,, 2014), this problem is well-documented (BlackJax,, 2019; Sountsov et al.,, 2024; Radul et al.,, 2020). However this also affects various other algorithms, such as variants of slice sampling (Neal,, 2003; Murray et al.,, 2010; Cabezas and Nemeth,, 2023), delayed rejection methods (Mira et al.,, 2001; Modi et al.,, 2024) and unbiased Gibbs sampling (Qiu et al.,, 2019).
In this work, we show how to transform MCMC algorithms into an equivalent sampling algorithm that avoids these synchronization barriers when using vmap-style vectorization. In particular,
@@
    We develop a novel approach to transform MCMC algorithms into finite state machines (FSMs), that can avoid synchronisation barriers when vectorizing with tools like vmap.
@@
    We implement several popular MCMC algorithms as FSMs, including Elliptical Slice Sampling, HMC-NUTS, and Delayed-Rejection MH - demonstrating speed-ups of up to an order of magnitude in experiments.
@@ 2.2 Synchronization Problems with While Loops
Control flow (i.e., if/else, while, for, etc.) poses a challenge for vectorization, because different batch members may require a different sequence of instructions. vmap solves this by executing all instructions for all batch members, and masking out the irrelevant computations. A consequence of this is that if sample contains a while loop, then vmap(sample) will execute the body of this loop for all chains until all termination conditions are met. Until then, no further instruction can be executed. As a result, if there is high variation in the number of loop iterations across chains, running Algorithm 1 with vmap(sample) introduces a synchronization barrier across all chains: at every iteration, each chain has to wait for the slowest sample call.
@@
Formalizing The Problem. Here we formalize the problem through a series of short derivations. These will be made precise in Section 4. Suppose we run a vmap’ed version of Algorithm 1 using a sample function that has a while loop. Let $N_{i,j}$ denote the number of iterations required by the $j$th chain to obtain its $i$th sample. If the while loop has a variable length, $N_{i,j}$ is a random number. Due to the synchronization problem described above, the time taken to run vmap(sample) at iteration $i$ is approximately proportional to the largest $N_{i,j}$ out of the $m$ chains, $\max_{j\leq m}N_{i,j}$. The total runtime after $n$ samples, $C_{0}(n)$, is then approximately proportional to,
@@
The key difference is that the maximum is now outside the sum. This reflects the fact that when running independently, we only have to wait at the end for the slowest chain to collect its $n$ samples, rather than waiting at every iteration. Clearly $C_{*}(n)\leq C_{0}(n)$. Significant speed-ups are obtainable by de-synchronizing the chains when $C_{*}(n)\ll C_{0}(n)$. If each $N_{i,j}$ converges in distribution to some $\mathbb{P}_{N}$ as we draw more samples and an appropriate law of large numbers holds, we can expect for large enough $n$ that:
[…]
where $(N_{\infty,1},...,N_{\infty,m})\overset{iid}{\sim}\mathbb{P}_{N}$ Therefore, de-synchronizing the chains should lead to large speedups if
@@
In this work, we will see that Equation 13 holds in various situations.
@@
On Figure 1 we display results when implementing this algorithm in JAX to sample from the hyperparameter posterior of a Gaussian process implemented on a regression task using a real dataset from the UCI repository (details are in Section 7). On the LHS we can see the distribution of the number of while loop iterations (i.e. slice shrinks) needed to generate a sample. While the average is $\sim$6, the average of the maximum across 1024 chains is $>$18, which implies $\mathbb{E}\max_{j\leq 1024}N_{n,j}\approx 3\mathbb{E}N_{n,1}$. On the RHS, we can see that the differences across the chains do balance out as more samples are drawn, which suggests a certain law of large numbers does hold here. In particular, after just 100 samples the distribution of the average number of iterations per chain is contained in the interval $[5,8]$, and after 10,000 samples this shrinks to $[6.2,6.4]$. This implies that if we could vectorize the algorithm without incurring synchronization barriers, we could improve the Effective Sample Size per Second (ESS/Sec) by up to 3-fold. We will see that for other algorithms (e.g., HMC-NUTS and delayed rejection) the potential speed-ups are much larger than this.
@@ 3   Finite State Machines for MCMC
In this section we present an implementation approach for MCMC algorithms with a given sample function, which avoids the above synchronization problems when vectorizing with vmap. The basic idea is to break down sample into a series of smaller ‘steps’ which avoid using iterative control flow like while loops and have minimal variance in execution time. We will then define a runtime procedure that allows chains to progress through their own step sequences in de-synchronized fashion. To do this in a principled manner, we use the framework of finite state machines (FSMs). Formally, an FSM is a 5-tuple $(\mathcal{S},\mathcal{Z},\delta,\mathcal{B},\mathcal{F})$, where $\mathcal{S}$ is a finite set of states, $\mathcal{Z}$ is a (finite) input set, $\delta:S\times\mathcal{Z}\to S$ is a transition function, $\mathcal{B}\in S$ is an initial state, and $\mathcal{F}$ is the set of ‘final’ states (Hopcroft et al.,, 2001). FSMs are useful tools which are typically used to represent algorithms that run on machines with finite storage. Below we show how to represent different sample functions as an FSM.
@@
At a high-level, we construct the FSM of an MCMC algorithm as follows. The states $\mathcal{S}=(S_{1},...,S_{K})$ are chosen as functions which execute contiguous code blocks of the algorithm. The boundaries of each code block are delineated by the start and end of any while loops. For example, $S_{1}$ executes all lines of code before the first while loop, $S_{2}$ executes all lines of code from the beginning of the first while loop body to either the start of the next while loop or the end of the current while loop, and so on. The inputs $z\in\mathcal{Z}$ are all variables taken as input to each code block. (e.g. current sample $\bm{x}$, proposal $\bm{x}^{\prime}$, log-likelihood $\log p(\bm{x})$) [^2] Whilst we may have that (e.g.) the sample $\bm{x}\in\mathbb{R}^{d}$, which violates finiteness of $\mathcal{Z}$, in practice an MCMC algorithm works using a finite subset of $\mathbb{R}^{d}$ (e.g. 32-bit floating point precision)., whilst the transition function $\delta$ selects the next code block to run according to the relevant loop condition and the received output $z^{\prime}$ after executing the current block (e.g. if a while loop starts after the block $S_{1}$ executes, $\delta$ will check this loop’s condition using the output $z^{\prime}=S_{1}(z)$). Below we make this construction procedure more precise for three kinds of sample functions which cover all MCMC algorithms considered in the present work: (i) functions with a single while loop, (ii) functions with two sequential while loops, and (iii) functions with two nested while loops. Additional details on the construction process are in Appendix B.
@@
Two sequential while loops. We break down sample into two blocks: $B_{1}$ contains all the code up to the second while-loop. $B_{2}$ contains the remaining code. In this case, $B_{1}$ and $B_{2}$ are now single while-loop programs, and thus can both be represented by FSMs $F_{1}$ and $F_{2}$ using the above rule. The FSM representation of sample can then be obtained by “stitching together” $F_{1}$ and $F_{2}$. The full construction process is in Appendix B. The resulting FSM is provided for the case of the Slice Sampler (Figure 3, top-right panel), which contains two [^3] We note that for 1D problems, the slice expansion loop can be broken into two loops—one for the upper bound of the interval, and one for the lower bound. while loops—one for expanding the slice, and one for contracting the slice.
Two nested while-loops In the case of two nested while loops, we break down sample into $B_{1},B_{2},B_{3}$, where $B_{1}$ (resp. $B_{3}$) is the code before (resp. after) the outer while loop and $B_{2}$ is the outer while-loop body. As $B_{2}$ is a single-while loop program, it admits its own FSM $F_{i}$. Building the final FSM of sample then informally consists in obtaining a first, “coarse” FSM by treating $B_{2}$ as opaque, and then refining it by replacing $B_{2}$ with its own FSM. The full construction process is given in Appendix B. The resulting FSM is provided for the case of NUTS (Figure 3, bottom-right panel), which—in its iterative form—uses the outer while loop to determine whether to keep expanding a Hamiltonian trajectory, and the inner while loop to determine whether to keep integrating along the current trajectory.
@@ 3.2 Defining the FSM runtime
Going forward, for convenience we assume the transition function $\delta$ takes in and returns the label $k$ of each block $S_{k}$, rather than $S_{k}$ itself. Now, given a constructed FSM, in Algorithm 2 we define a function step(k,z) which when executed performs a single transition along an edge in the FSM graph. For reasons that will be clear shortly, we augment this function with a flag that indicates when the final block is run.
@@
Note that if we start from some input $(k,z)=(0,z_{0})$, by calling step repeatedly we will transition through a sequence of blocks of sample, until we eventually reach the terminal state, at which point a sample is obtained (as indicated by isSample=True). We can use this function to draw $n$ samples, by (1) adding a transition from the terminal state $\mathcal{F}=S_{K}$ back to the initial state $\mathcal{B}=S_{0}$, and (2) defining a wrapper function which iteratively calls step until isSample=True is obtained $n$-times (see Algorithm 3).
@@
Both Algorithm 3 and Algorithm 1 draw $n$ samples from sample and can be easily vectorized with vmap. In the case of Algorithm 3, we just call vmap on step and modify the outer loop to terminate when all chains have collected $n$ samples each (see Algorithm 7 in Appendix B). The crucial difference is that: (i) by changing the definition of a ‘step’ to allow different chains to execute different code blocks, Algorithm 3 enables them to progress their own independent block sequences, and (ii) by essentially moving the while loop to the outer layer, Algorithm 3 only requires the chains to re-synchronize after $n$ samples.
@@ 4   Time Complexity Analysis of FSM-MCMC
One limitation with our FSM design is the step function relies on a switch to determine which block to run. When vectorized with vmap or an equivalent transformation, all branches are evaluated for all chains, with irrelevant results discarded. This means the cost of a single call of step is the cost of running all state functions $S_{1},...,S_{K}$. To obtain a speedup, the FSM must therefore sufficiently decrease the expected number of steps to obtain $n$ samples from each chain. In this section we quantitatively derive conditions under which this occurs in the simplified setting of an MCMC algorithm with a single while loop. This enables us to subequently optimize the design of the FSM in Section 5.
@@
To this end, consider a sample function with a single while loop, and associated FSM with $K$ states, where $S_{k}$ (for some $k\in[K]$) executes the body of this loop. Note our approach yields $K\leq 3$ when there is one while loop, but we relax this here to analyze the effect of the number of blocks on performance. We assume the cost of executing $S_{1},..,S_{K}$ is $c_{1}(m),...c_{K}(m)$ respectively when using vectorized Algorithm 1 and $\alpha\sum_{k=1}^{K}c_{k}(m)$ when using vectorized Algorithm 3, where $\alpha\in[\max_{k\in[K]}c_{k}(m)/\sum_{k=1}^{K}c_{k}(m),1]$. The variable $\alpha$ reflects the cost of using multiple branches in step, which is $1$ using Algorithm 2, but can be decreased by refining step, as we show later. The dependence of each cost $c_{i}(m)$ on $m$ reflects the ability of the GPU to efficiently parallelize a call to code-block $S_{i}$ for $m$ chains. Under these assumptions, the average cost per sample for $m$ chains after $n$ samples each, is expressed for the standard design $C_{0}(n,m)$ and FSM design $C_{F}(n,m)$ as
$C_{0}(m,n)$ $\;=\;A_{0}(m)+B_{0}(m)\frac{1}{n}\sum_{i=1}^{n}\max_{j\in[m]}\left(N_{i,j}\right)$ (14)
$C_{F}(m,n)$ $\;=\;A_{F}(m)+B_{F}(m)\max_{j\in[m]}\left(\frac{1}{n}\sum_{i=1}^{n}N_{i,j}\right)$ (15)
Here $A_{0}(m)=c_{\neg k}(m)$ and $A_{F}(m)=\alpha(K-1)(c_{\neg k}(m)+c_{k}(m))$ are the costs of calling all blocks except $S_{k}$ (and $c_{\neg k}(m)=\sum_{j\neq k}c_{j}(m)$); $B_{0}(m)=c_{k}(m)$ and $B_{F}(m)=\alpha(c_{k}(m)+c_{\neg k}(m))$ are the costs of running the iterative block $S_{k}$; and $N_{i,j}$ is now the number of calls of $S_{k}$ needed to produce sample $i$ for chain $j$ (i.e. $\bm{X}_{i,j}$) [^4] We use capital letters here as our analysis in this section treats the samples as random variables.. If the joint sequence $(\bm{X}_{i,j},N_{i,j})_{i\geq 1}$ is a convergent Markov chain, we can state the following concentration result, which formally justifies our derivations leading to (9) and (12).
Theorem 4.1.
@@
Let $N_{i,j}\in[0,B],\bm{X}_{i,j}\in\mathcal{X}\subset\mathbb{R}^{d}$, $(\bm{X}_{i,j},N_{i,j})_{i\geq 1}$ be a Markov Chain with stationary distribution $\pi$. Then with probability $1-\delta$ we have the inequalities
$|C_{0}(m,n)-A_{0}(m)-B_{0}(m)\mathbb{E}_{\pi}\max_{j\in[m]}N_{j}|\;\leq\;MB_{0}(m)n^{-\frac{1}{2}}\ln(2/\delta)$ (16)
$|C_{F}(m,n)-A_{F}(m)-B_{F}(m)\mathbb{E}_{\pi}N_{1}|\;\leq\;MB_{F}(m)n^{-\frac{1}{2}}\ln(2m/\delta)$ (17)
Where $(N_{1},..,N_{m})\overset{iid}{\sim}\mathbb{\pi}_{N}$ and $M>0$.
@@
The result essentially says that as ${n\to\infty}$, the cost per sample of the standard MCMC design will converge to the expected time for the slowest chain to draw a sample (i.e. $A_{0}(m)-B_{0}(m)\mathbb{E}_{\pi}\max_{j\in[m]}N_{j}$), whilst the FSM design will converge to the expected time for a single chain to draw a sample, scaled by an additional cost of control-flow in step (i.e. $A_{F}(m)-B_{F}(m)\mathbb{E}_{\pi}N_{1}|$). Both convergence rates are $\mathcal{O}(n^{-\frac{1}{2}})$.
Remark 4.2.
The Markov chain assumption is satisfied whenever $N_{i,j}$ depends (only) on the geometry of the distribution at $\bm{X}_{i-1,j}$. Its convergence can be induced by irreducibility and aperiodicity, which we expect will typically hold under irreducibility and aperiodicity of the marginal Markov chain $(\bm{X}_{i,j})_{i>1}$.
4.1 The long-run relative cost of the FSM

### v2

@@ 1 Introduction
Automatic vectorization is the act of transforming a function to handle batches of inputs without user intervention. Implementations of automatic vectorization algorithms are now available in many mainstream scientific computing libraries, and have dramatically simplified the task of running multiple instances of a single algorithm concurrently. They are routinely used to train neural networks (Flax,, 2023) and in other scientific applications, e.g., Schoenholz and Cubuk, (2021); Oktay et al., (2023); Pfau et al., (2020).
This paper focuses on the use of automatic vectorization for Markov chain Monte Carlo (MCMC) algorithms. Tools like JAX’s vmap provide a convenient way to run multiple MCMC chains in parallel: one can simply write single-chain MCMC code, and call vmap to turn it into vectorized, multi-chain code that can run in parallel on the same processor. Many state-of-the-art MCMC libraries have consequently adopted machine learning frameworks with automatic vectorization tools as their backend.
One limitation with automatic vectorization tools is how they handle control flow. Since all instructions must be executed in lock-step, if the algorithm has a while loop, all chains must wait until the last chain has finished its iterations. This can lead to large inefficiencies for MCMC algorithms that generate each sample using variable-length while loops. Roughly speaking, if vectorization executes, say, 100 chains in parallel, all but one finish after at most 10 steps, and the remaining chain runs for 1000 steps, then about 99% of the GPU capacity assigned to vmap is wasted (and our simulations show the effect can indeed be this drastic). For the No-U-Turn Sampler (HMC-NUTS) (Hoffman et al.,, 2014), this problem is well-documented (BlackJax,, 2019; Sountsov et al.,, 2024; Radul et al.,, 2020). It also affects various other algorithms, such as variants of slice sampling (Neal,, 2003; Murray et al.,, 2010; Cabezas and Nemeth,, 2023), delayed rejection methods (Mira et al.,, 2001; Modi et al.,, 2024) and unbiased Gibbs sampling (Qiu et al.,, 2019).
In this work, we show how to transform MCMC algorithms into equivalent samplers that avoid these synchronization barriers when using vmap-style vectorization. In particular,
@@
    We develop a novel approach to transform MCMC algorithms into finite state machines (FSMs), that can avoid synchronization barriers with vmap.
@@
    We implement popular MCMC algorithms as FSMs, including Elliptical Slice Sampling, HMC-NUTS, and Delayed-Rejection MH - demonstrating speed-ups of up to an order of magnitude in experiments.
@@ 2.2 Synchronization Problems with While Loops
Control flow (i.e., if/else, while, for, etc.) poses a challenge for vectorization, because different batch members may require a different sequence of instructions. vmap solves this by executing all instructions for all batch members, and masking out irrelevant computations. As a consequence, if sample contains a while loop, then vmap(sample) will execute the body of this loop for all chains until all termination conditions are met. Until then, no further instruction can be executed. As a result, if there is high variation in the number of loop iterations across chains, running Algorithm 1 with vmap(sample) introduces a synchronization barrier across all chains: at every iteration, each chain has to wait for the slowest sample call.
@@
Formalizing The Problem. Here we formalize the problem through a series of short derivations. These will be made precise in Section 4. Suppose we run a vmap’ed version of Algorithm 1 using a sample function that has a while loop. Let $N_{i,j}$ denote the number of iterations required by the $j$th chain to obtain its $i$th sample. If the while loop has a variable length, $N_{i,j}$ is a random number. Due to the synchronization problem described above, the time taken to run vmap(sample) at iteration $i$ is approximately proportional to the largest $N_{i,j}$ out of the $m$ chains, $\max_{j\leq m}N_{i,j}$. If we ignore overheads, the total runtime after $n$ samples, $C_{0}(n)$, is then roughly proportional to
@@
The key difference is that the maximum is now outside the sum. This reflects the fact that when running independently, we only have to wait at the end for the slowest chain to collect its $n$ samples, rather than waiting at every iteration. Clearly $C_{*}(n)\leq C_{0}(n)$. Significant speed-ups are obtainable by de-synchronizing the chains when $C_{*}(n)\ll C_{0}(n)$. If each $N_{i,j}$ converges in distribution to some $\mathbb{P}_{N}$ as we draw more samples, and an appropriate central limit theorem holds (so that the sums in (3) and (6) behave like scaled means), we can expect for large enough $n$ that:
$C_{0}(n)$ $\;=\;\mathcal{O}_{p}(n\,\mathbb{E}\max_{j\leq m}\,N_{\infty,j})$ (7)
$C_{*}(n)$ $\;=\;\mathcal{O}_{p}(\max_{j\leq m}\,n\,\mathbb{E}N_{\infty,j})\;=\;\mathcal{O}_{p}(n\,\mathbb{E}N_{\infty,1})$ (8)
where the last equality assumes $(N_{\infty,1},...,N_{\infty,m})\overset{iid}{\sim}\mathbb{P}_{N}$. De-synchronizing the chains will result in large speedups if
@@
We will see that Equation 9 holds in various situations.
@@
On Figure 1 we display results when implementing this algorithm in JAX to sample from the hyperparameter posterior of a Gaussian process implemented on a regression task using a real dataset from the UCI repository (details are in Section 7). On the LHS we can see the distribution of the number of while loop iterations (i.e. slice shrinks) needed to generate a sample. While the average is $\sim$6, the average of the maximum across 1024 chains is $>$18, which implies $\mathbb{E}\max_{j\leq 1024}N_{n,j}\approx 3\mathbb{E}N_{n,1}$. On the RHS, we can see that the differences across the chains do balance out as more samples are drawn, which suggests a certain central limit theorem may hold. In particular, after just 100 samples the distribution of the average number of iterations per chain is contained in the interval $[5,8]$, and after 10,000 samples this shrinks to $[6.2,6.4]$. This implies that if we could vectorize the algorithm without incurring synchronization barriers, we could improve the Effective Sample Size per Second (ESS/Sec) by up to 3-fold. We will see that for other algorithms (e.g., HMC-NUTS and delayed rejection) the potential speed-ups are much larger than this.
@@ 3 Finite State Machines for MCMC
In this section we present an approach to implementing MCMC algorithms, which avoids the above synchronization problems when vectorizing with vmap. The basic idea is to break down the transition kernel (sample) into a series of smaller ‘steps’ which avoid iterative control flow like while loops and have minimal variance in execution time. We then define a runtime procedure that allows chains to progress through their own step sequences in de-synchronized fashion. To do this in a principled manner, we adapt the framework of finite state machines (FSMs) (Hopcroft et al.,, 2001). An FSM is a 5-tuple $(\mathcal{S},\mathcal{Z},\delta,\mathcal{B},\mathcal{F})$, where $\mathcal{S}$ is a finite set of states, $\mathcal{Z}$ is a finite input set, $\delta:S\times\mathcal{Z}\to S$ is a transition function, $\mathcal{B}\in S$ is an initial state, and $\mathcal{F}$ is the set of final states. We deviate from the usual definition by allowing each state to be a function $S:\mathcal{Z}\to\mathcal{Z}$ that updates inputs dynamically [^2] This essentially results in a so-called ‘Finite State Transducer’ with a different runtime than typical.. Below we show how to represent different sample functions as an FSM.
@@
At a high-level, we construct the FSM of an MCMC algorithm as follows. The states $\mathcal{S}=(S_{1},...,S_{K})$ are chosen as functions $\mathcal{Z}\to\mathcal{Z}$ which execute contiguous code blocks of the algorithm. The boundaries of each code block are delineated by the start and end of any while loops. For example, $S_{1}$ executes all lines of code before the first while loop, $S_{2}$ executes all lines of code from the beginning of the first while loop body to either the start of the next while loop or the end of the current while loop, and so on. The inputs $z\in\mathcal{Z}$ are all variables used in each code block. (e.g. current sample $\bm{x}$, proposal $\bm{x}^{\prime}$, log-likelihood $\log p(\bm{x})$) [^3] Typically $\bm{x}\in\mathbb{R}^{d}$ so $\mathcal{Z}$ is not finite. However, in practice all samples lie in a finite subset of $\mathbb{R}^{d}$ (e.g. 32-bit floating points)., whilst the transition function $\delta$ selects the next code block to run according to the relevant loop condition and the received output $z^{\prime}$ after executing the current block (e.g. if a while loop starts after the block $S_{1}$ executes, $\delta$ will check this loop’s condition using the output $z^{\prime}=S_{1}(z)$). Below we make this construction procedure more precise for three kinds of sample functions which cover all MCMC algorithms considered in the present work: (i) functions with a single while loop, (ii) functions with two sequential while loops, and (iii) functions with two nested while loops. An automated construction process for more general programs using a toy language is given in Appendix B.
@@
Two Sequential While Loops. We break down sample into two blocks: $B_{1}$ contains all the code up to the second while loop. $B_{2}$ contains the remaining code. In this case, $B_{1}$ and $B_{2}$ are now single while loop programs, and thus can both be represented by FSMs $F_{1}$ and $F_{2}$ using the above rule. The FSM representation of sample can then be obtained by “stitching together” $F_{1}$ and $F_{2}$. The full construction process is in Appendix B. The resulting FSM is provided for the case of the Slice Sampler (Figure 3, top-right panel), which contains two [^4] For 1D problems, the slice expansion loop can be broken into two loops (for the upper and lower bound of the interval). while loops: one for expanding the slice, and one for contracting the slice.
Two Nested While Loops. In the case of two nested while loops, we break down sample into $B_{1},B_{2},B_{3}$, where $B_{1}$ (resp. $B_{3}$) is the code before (resp. after) the outer while loop and $B_{2}$ is the outer while loop body. As $B_{2}$ is a single-while loop program, it admits its own FSM $F_{i}$. Building the final FSM of sample then informally consists in obtaining a first, “coarse” FSM by treating $B_{2}$ as opaque, and then refining it by replacing $B_{2}$ with its own FSM. The full construction process is given in Appendix B. The resulting FSM is provided for the case of NUTS (Figure 3, bottom-right panel), which—in its iterative form—uses the outer while loop to determine whether to keep expanding a Hamiltonian trajectory, and the inner while loop to determine whether to keep integrating along the current trajectory.
@@ 3.2 Defining the FSM Runtime
Going forward, for convenience we assume the transition function $\delta$ takes in and returns the label $k$ of each block $S_{k}$, rather than $S_{k}$ itself. Now, given a constructed FSM, in Algorithm 2 we define a function step, which when executed performs a single transition along an edge in the FSM graph. For reasons made clear shortly, we augment step to return a flag indicating when the final block is run.
@@
Here $switch|$ uses the value of $k$ to determine which branch to run. If we start from some input $(k,z)=(0,z_{0})$ and call step repeatedly, we transition through a sequence of blocks of sample, until we eventually reach the terminal state. At this point, a sample is obtained (as indicated by isSample=True). We can use this function to draw $n$ samples, by (i) adding a transition from the terminal state $\mathcal{F}=S_{K}$ back to the initial state $\mathcal{B}=S_{0}$, and (ii) defining a wrapper function which iteratively calls step until isSample=True is obtained $n$-times (see Algorithm 3).
@@
Both Algorithm 3 and Algorithm 1 call sample $n$-times and can be easily vectorized with vmap. For Algorithm 3, we just call vmap on step and modify the outer loop to terminate when all chains have collected $n$ samples each (see Algorithm 7 in Appendix B). The crucial difference is: (i) by defining a ‘step’ to be agnostic to which block each chain is currently on, Algorithm 3 enables chains to simultaneously progress their independent block sequences, and (ii) by moving all while loops to the outer layer, Algorithm 3 only requires chains to synchronize after $n$ samples.
@@ 4 Time Complexity Analysis of FSM-MCMC
One limitation with our FSM design is the step function relies on a switch to determine which block to run. When vectorized with vmap or an equivalent transformation, all branches in switch are evaluated for all chains, with irrelevant results discarded. This means the cost of a single call of step is the cost of running all state functions $S_{1},...,S_{K}$. To obtain a speedup, the FSM must therefore sufficiently decrease the expected number of steps to obtain $n$ samples from each chain. In this section we quantitatively derive conditions under which this occurs in the simplified setting of an MCMC algorithm with a single while loop. This enables us to subsequently optimize the FSM design in Section 5.
4.1 Long Run Costs of FSM and Non-FSM MCMC
@@
To this end, consider a sample function with a single while loop. Suppose it is broken down into $K$ blocks $\{S_{1},...S_{K}\}$, by our FSM construction procedure, where $S_{k}$ (for some $k\in[K]$) executes the body of the loop. Note our procedure yields $K\leq 3$ for one while loop, but we relax this here to analyze the effect of the number of blocks on performance. A single call of vmap(sample) executes $S_{l}$ once if $l\neq k$, and $\max_{j}N_{j}$-times if $l=k$, where $N_{j}=$ loop iterations for chain $j$. We assume the cost of executing each block $S_{l}$ with vmap is $c_{l}(m)$ for $m$ chains (here the dependence on $m$ reflects underlying GPU vectorization efficiency). The average cost per sample of Algorithm 1 after $n$ samples, when vectorized for $m$ chains, is then
$C_{0}(m,n)$ $\;:=\;A_{0}(m)+B_{0}(m)\frac{1}{n}\sum_{i=1}^{n}\max_{j\in[m]}\,N_{i,j}$ (10)
where $A_{0}(m):=\sum_{j\neq k}c_{j}(m)$, $B_{0}(m)=c_{k}(m)$, and $N_{i,j}$ = # calls of the loop body $S_{k}$ for chain $j$ and sample $i$. The $\max$ reflects the synchronization barrier across chains.
In this case, for the FSM (vmap’ed Algorithm 3), the cost of calling each block $S_{l}$ is $\sum_{j=1}^{K}c_{j}(m)$, since vmap(step) executes all switch branches for all chains. We will assume this cost can be scaled by some $\alpha\in[\max_{k\in[K]}c_{k}(m)/\sum_{k=1}^{K}c_{k}(m),1]$, since we will later redesign step to reduce $\alpha<1$. This gives an average cost per sample for $m$ chains after $n$ samples of,
$C_{F}(m,n)$ $\;=\;A_{F}(m)+B_{F}(m)\max_{j\in[m]}\,\frac{1}{n}\sum_{i=1}^{n}N_{i,j}$ (11)
where $A_{F}(m)=\alpha(K-1)(c_{\neg k}(m)+c_{k}(m))$, $B_{F}(m)=\alpha(c_{k}(m)+c_{\neg k}(m))$, and $c_{\neg k}(m):=\sum_{j\neq k}c_{j}(m)$.
Let $\bm{X}_{i,j}$ be the $i^{th}$ random sample obtained by chain $j$. If the joint sequence $(\bm{X}_{i,j},N_{i,j})_{i\geq 1}$ is an appropriate Markov chain, we have the following concentration result for $C_{0},C_{F}$, which formalizes our derivations in Section 2.2.
Theorem 4.1 (Long Run Costs).
@@
Let $N_{i,j}\in[0,B],\bm{X}_{i,j}\in\mathcal{X}\subseteq\mathbb{R}^{d}$, $(\bm{X}_{i,j},N_{i,j})_{i\geq 1}$ be a Markov Chain with stationary distribution $\pi$ with absolute spectral gap $1-\lambda\in[0,1)$. Then with probability $1-\delta$ we have the inequalities:
$|C_{0}(m,n)-C_{0}(m)|\;$ $\leq\;MB_{0}(m)n^{-\frac{1}{2}}\ln(2/\delta)$ (12)
$|C_{F}(m,n)-C_{F}(m)|\;$ $\leq\;MB_{F}(m)n^{-\frac{1}{2}}\ln(2m/\delta)$ (13)
Where, for some $(N_{1},..,N_{m})\overset{iid}{\sim}\mathbb{\pi}_{N}$, the long run costs are
$C_{0}(m)$ $:=A_{0}(m)+B_{0}(m)\mathbb{E}_{\pi}\max_{j\in[m]}N_{j}$ (14)
$C_{F}(m)$ $:=A_{F}(m)+B_{F}(m)\mathbb{E}_{\pi}N_{1}$ (15)
@@
The result says that as ${n\to\infty}$, the cost per sample of the standard MCMC design converges to the expected time for the slowest chain to draw a sample, whilst the FSM design converges to the expected time for a single chain to draw a sample, scaled by the additional cost of control-flow in step (i.e. $A_{F}(m)$ and $B_{F}(m)$). Both rates are $\mathcal{O}(n^{-\frac{1}{2}})$.
4.2 The Relative Long-Run Cost of the FSM